OpenStudy (anonymous):

Find f(5) for this sequence: (I need the answer/how you got it so I can use it as a example for other problems) f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

f(5) = f(1) + f(2) + f(5-1) ? if so, f(5) = 2 + 3 + f(4) f(4) = f(1) + f(2) + f(4-1) = 2 + 3 + f(3) f(3) = f(1) + f(2) + f(3-1) = 2 + 3 + 3 therefore f(5) = 2 + 3 + (2 + 3 + (2 + 3 + 3)) = 5 + 5 + 8 = 18 (assuming it's a recursive function)